# fairmetrics: An R package for group fairness evaluation

### v2

@@ 4.1 Independence
  • Conditional Statistical Parity: Restricts the comparison of positive prediction rates to a specific subgroup (e.g., within a hospital unit or age bracket), offering a more context-specific fairness assessment.
@@ 5 Evaluating Fairness Criteria
\$performance
                                    Metric GroupFemale GroupMale
1                 Positive Prediction Rate        0.18      0.09
2                 Positive Prediction Rate        0.35      0.24
3                      False Negative Rate        0.36      0.59
4                      False Positive Rate        0.08      0.03
5                     Avg. Predicted Prob.        0.47      0.37
6                     Avg. Predicted Prob.        0.16      0.11
7                Positive Predictive Value        0.62      0.67
8                Negative Predictive Value        0.93      0.91
9                              Brier Score        0.09      0.08
10                                Accuracy        0.87      0.89
11 (False Negative)/(False Positive) Ratio        0.93      2.89
\$fairness
                           Metric Difference    95% Diff CI Ratio 95% Ratio CI
1              Statistical Parity       0.08   [0.040.12]  2.00 [1.412.83]
2  Conditional Statistical Parity       0.10   [0.020.18]  1.43 [1.051.96]
3               Equal Opportunity      -0.21 [-0.36-0.06]  0.64  [0.450.9]
4             Predictive Equality       0.04   [0.010.07]  2.33  [1.134.8]
5      Balance for Positive Class       0.09   [0.040.14]  1.24 [1.091.42]
6      Balance for Negative Class       0.05   [0.030.07]  1.50 [1.291.75]
7      Positive Predictive Parity      -0.06  [-0.230.11]  0.92 [0.711.18]
8      Negative Predictive Parity       0.01  [-0.150.17]  1.01 [0.791.29]
9              Brier Score Parity       0.01  [-0.010.03]  1.12 [0.891.43]
10        Overall Accuracy Parity      -0.01  [-0.050.03]  0.99 [0.951.03]
11             Treatment Equality      -2.28  [-4.720.16]  0.33 [0.150.73]

### v3

@@ 5 Evaluating Fairness Criteria
           Fairness Assesment                                  Metric
1          Statistical Parity                Positive Prediction Rate
2           Equal Opportunity                     False Negative Rate
3         Predictive Equality                     False Positive Rate
4  Balance for Positive Class           Avg. Predicted Positive Prob.
5  Balance for Negative Class           Avg. Predicted Negative Prob.
6  Positive Predictive Parity               Positive Predictive Value
7  Negative Predictive Parity               Negative Predictive Value
8          Brier Score Parity                             Brier Score
9     Overall Accuracy Parity                                Accuracy
10         Treatment Equality (False Negative)/(False Positive) Ratio
   GroupFemale GroupMale Difference    95% Diff CI Ratio 95% Ratio CI
1         0.17      0.08       0.09   [0.050.13]  2.12 [1.493.04]
2         0.38      0.62      -0.24 [-0.39-0.09]  0.61 [0.440.86]
3         0.08      0.03       0.05   [0.020.08]  2.67  [1.45.08]
4         0.46      0.37       0.09   [0.040.14]  1.24 [1.091.42]
5         0.15      0.10       0.05   [0.030.07]  1.50 [1.291.74]
6         0.62      0.66      -0.04  [-0.210.13]  0.94 [0.721.22]
7         0.92      0.90       0.02  [-0.020.06]  1.02 [0.981.07]
8         0.09      0.08       0.01  [-0.010.03]  1.12 [0.891.43]
9         0.87      0.88      -0.01  [-0.050.03]  0.99 [0.941.04]
10        1.03      3.24      -2.21 [-4.38-0.04]  0.32 [0.150.68]